# SIC-Free Rate-Splitting Multiple Access: Constellation-Constrained Optimization and Application to Large-Scale Systems

### v1

@@ III-B Achievable Rate of RSMA
$+\frac{1}{|\mathcal{X}_{\mathcal{K}}|}\sum_{m=1}^{|\mathcal{X}_{\mathcal{K}}|}\log_{2}\sum_{l=1}^{|\mathcal{X}_{\mathcal{K}}|}\exp\Big{(}-\frac{|\mathbf{d}_{m,l}(\sum_{a\in\mathcal{K}}\mathbf{h}^{H}\mathbf{p}_{a}\mathcal{X}_{a})|^{2}}{2\sigma^{2}}\Big{)}\Big{\}}$ (14)
@@
$+\frac{1}{|\mathcal{X}_{\mathcal{K}/k}|}\sum_{m=1}^{|\mathcal{X}_{\mathcal{K}/k}|}\log_{2}\sum_{l=1}^{|\mathcal{X}_{\mathcal{K}/k}|}\exp\Big{(}-\frac{|\mathbf{d}_{m,l}(\sum_{a\in\mathcal{K}/k}\mathbf{h}^{H}\mathbf{p}_{a}\mathcal{X}_{a})|^{2}}{2\sigma^{2}}\Big{)}\Big{\}}$ (15)
$R_{\text{p},k}^{\text{non-SIC, approx.}}=$ $\log_{2}|\mathcal{X}_{k}|-\frac{1}{|\mathcal{X}_{\text{c}}\times\mathcal{X}_{\mathcal{K}}|}\sum_{m=1}^{|\mathcal{X}_{\text{c}}\times\mathcal{X}_{\mathcal{K}}|}\log_{2}\sum_{l=1}^{|\mathcal{X}_{\text{c}}\times\mathcal{X}_{\mathcal{K}}|}\exp\Big{(}-\frac{|\mathbf{d}_{m,l}(\mathbf{h}^{H}\mathbf{p}_{\text{c}}\mathcal{X}_{\text{c}}+\sum_{a\in\mathcal{K}}\mathbf{h}^{H}\mathbf{p}_{a}\mathcal{X}_{a})|^{2}}{2\sigma^{2}}\Big{)}$
@@
Let $C_{k}$ denote the portion of the common stream rate that is allocated to user-$k$. Inherently, $C_{k}\geq 0$, $\forall k$, and $\sum_{k}C_{k}=R_{\text{c}}$. The rate experienced by user-$k$ is given by [^3] Hereinafter, the superscript ”SIC/non-SIC” indicates that the expression depends on whether SIC receivers or non-SIC receivers are used.
$R_{k}^{\text{SIC/non-SIC}}=C_{k}+R_{\text{p},k}^{\text{SIC/non-SIC}}.$ (17)
@@ IV Precoding Optimization
In this section, we consider WSR and MMF maximization problems for RSMA with SIC and non-SIC receivers, and propose precoding optimization algorithms to address them.
@@ IV-A Weighted sum-rate optimization
$\begin{split}\mathcal{P}_{1}:\;\;\underset{\mathbf{P},\mathbf{c}}{\max}\>\>&\sum_{k\in\mathcal{K}}u_{k}R_{k}^{\text{SIC/non-SIC}}\\ \text{s.t.}\>\>\>\;&\|\mathbf{P}\|_{\text{F}}^{2}\leq P_{\text{T}}\\ &\mathbf{c}^{T}\mathbf{1}=R_{\text{c}}\\ &\mathbf{c}\succcurlyeq\mathbf{0},\end{split}$ (18)
@@
As a direct consequence of Proposition 1, one solution of $\mathbf{c}$ can already be determined given $u_{k}$, $k\in\mathcal{K}$. As a result, $\mathbf{c}$ can be removed from the design variables, and $\mathcal{P}_{1}$ is simplified into
$\begin{split}\mathcal{P}_{2}:\;\;\underset{\mathbf{P}}{\max}\>\>&u_{k^{\prime}}R_{\text{c}}+\sum_{k\in\mathcal{K}}u_{k^{\prime}}R_{\text{p},k}^{\text{SIC/non-SIC}}\\ \text{s.t.}\>\>\>\;&\|\mathbf{P}\|_{\text{F}}^{2}\leq P_{\text{T}},\end{split}$ (19)
@@
The gradient of (6) is given by
$\begin{split}&\nabla_{\mathbf{P}}H(\mathbf{s}|y,\mathbf{h}^{H}\mathbf{P})^{\text{approx.}}\\ &=\frac{\log_{2}e}{|\mathcal{X}|}\sum_{m=1}^{|\mathcal{X}|}\frac{1}{\sum_{l=1}^{|\mathcal{X}|}u(\mathbf{P})}\sum_{l=1}^{|\mathcal{X}|}u(\mathbf{P})\left(-\frac{\mathbf{hh}^{H}\mathbf{P}\mathbf{d}_{m,l}\mathbf{d}_{m,l}^{H}}{2\sigma^{2}}\right),\end{split}$ (20)
@@
$u(\mathbf{P})=\sum_{l=1}^{|\mathcal{X}|}\exp\Big{(}-\frac{|\mathbf{h}^{H}\mathbf{P}\mathbf{d}_{m,l}|^{2}}{2\sigma^{2}}\Big{)}.$ (21)
Based on (20), $\nabla_{\mathbf{P}}R_{\text{c},k}$, $\nabla_{\mathbf{P}}R_{k}^{\text{SIC}}$ and $\nabla_{\mathbf{P}}R_{k}^{\text{non-SIC}}$ can be obtained trivially. One subgradient of the weighted sum-rate can be obtained as
$g(R_{\text{WSR}}^{\text{SIC/non-SIC}})=\max_{k}u_{k}\nabla_{\mathbf{P}}R_{\text{c},k^{\prime}}+\sum_{k\in\mathcal{K}}u_{k}\nabla_{\mathbf{P}}R_{k}^{\text{SIC/non-SIC}},$ (22)
where $k^{\prime}=\arg\min_{k\in\mathcal{K}}R_{\text{c},k}$.
It is easy to see that any solution of $\mathcal{P}_{2}$, $\mathbf{P}^{\star}$, must satisfy $\|\mathbf{P}^{\star}\|_{\text{F}}^{2}=P_{\text{T}}$. We therefore use the projection of $\mathbf{P}$ in $\{\mathbf{P}\big{|}\|\mathbf{P}\|_{\text{F}}^{2}=P_{\text{T}}\}$ as the updated solution after every iteration. The proposed Projected SubGradient Ascent (PSGA) algorithm is summarized in Algorithm 1444In Algorithm 1, depending on the assumption about receivers, $R_{\text{WSR}}=R_{\text{WSR}}^{\text{SIC}}\text{ or }R_{\text{WSR}}^{\text{non-SIC}}$.. For choosing step length, $t$, we use a modified backtracking line search specific for PSGA as described in Algorithm 2. Compared to the classical backtracking for gradient descent/ascent, the proposed backtracking method incorporates projection when evaluating the predicted objective value after updating, ensuring a non-decreasing sequence of objectives. The convergence of general projected gradient methods is discussed in [39], although the convergence rate of PSGA for non-convex problems is generally difficult to prove. Note that Algorithm 2 leads to non-decreasing variation to the objective function, and the objective function of (18) is upper bounded, therefore, Algorithm 1 is guaranteed to converge. The convergence is also numerically evaluated in Figure 5, where a monotonic increase in the objective function and the convergence can be observed.
@@ IV-B Max-min fairness optimization
$\begin{split}\mathcal{P}_{3}:\;\;\underset{\mathbf{P},\mathbf{c}}{\max}\>\>&\min_{k\in\mathcal{K}}\>\>R_{k}^{\text{SIC/non-SIC}}\\ \text{s.t.}\>\>\>\;&\|\mathbf{P}\|_{\text{F}}^{2}\leq P_{\text{T}}\\ &\mathbf{c}^{T}\mathbf{1}=R_{\text{c}}\\ &\mathbf{c}\succcurlyeq\mathbf{0}.\end{split}$ (23)
Note that $\mathcal{P}_{3}$ is also a non-convex problem and hence is difficult to solve. To develop efficient algorithms for $\mathcal{P}_{3}$, we decouple it into the following two sub-problems,
$\begin{split}\mathcal{P}_{4}:\;\;\underset{\mathbf{c}}{\text{max}}\>\>&\min_{k\in\mathcal{K}}\>\>R_{k}(\mathbf{c};\mathbf{P})\\ \text{s.t.}\>\>\>\;&\mathbf{c}^{T}\mathbf{1}=R_{\text{c}}\\ &\mathbf{c}\succcurlyeq\mathbf{0},\end{split}$ (24)
@@
where $R_{k}(\mathbf{c};\mathbf{P})$ indicates the case where $R_{k}^{\text{SIC/non-SIC}}$ is determined with $\mathbf{P}$ as a given parameter and $\mathbf{c}$ as a variable, whereas $R_{k}(\mathbf{P};\mathbf{c})$ indicates the case in which $\mathbf{c}$ is given and $\mathbf{P}$ is the variable. Hence, $\mathcal{P}_{4}$ optimizes $\mathbf{c}$ with a fixed $\mathbf{P}$ while $\mathcal{P}_{5}$ optimizes $\mathbf{P}$ with a given $\mathbf{c}$.
@@
4:     $\mathbf{d}:=[R_{\text{p},1},...,R_{\text{p},k^{\prime}}]^{T}$.
5:     $\eta:=\frac{R_{\text{c}}+\mathbf{1}^{T}\mathbf{d}}{k^{\prime}}$.
6:     $\mathbf{c}^{\prime}:=\eta\mathbf{1}-\mathbf{d}$.
@@
The gap between point-wise minimum and its log-sum-exp approximation is characterized by the following [40],
$00\leq\min_{1\leq i\leq L}f_{i}-\frac{1}{\gamma}\log\sum_{i}\exp(\gamma f_{i})\leq\frac{1}{\gamma}\log L,\;\;\;\gamma<$ 0 . (27)
Let $R_{\text{max-min}}$ denotes the objective function in (26). One subgradient of $R_{\text{max-min}}$ w.r.t. $\mathbf{P}$ is derived to be
$g(R_{\text{max-min}})=\frac{1}{\gamma}\frac{\sum_{k\in\mathcal{K}}\gamma\exp\big{(}\gamma R_{k}(\mathbf{P};\mathbf{c})\big{)}g\big{(}R_{k}(\mathbf{P};\mathbf{c})\big{)}}{\sum_{k\in\mathcal{K}}\exp\big{(}\gamma R_{k}(\mathbf{P};\mathbf{c})\big{)}},$ (28)
@@
$g(R_{k})=c_{k}\nabla_{\mathbf{P}}R_{\text{c},k^{\prime}}+\nabla_{\mathbf{P}}R_{\text{p},k},$ (29)
and $k^{\prime}=\arg\min_{k}R_{\text{c},k}$. Similar to $\mathcal{P}_{2}$, $\mathcal{P}_{6}$ can also be solved using PSGA. The solution of (23) is obtained by updating the solution for $\mathcal{P}_{4}$ and $\mathcal{P}_{6}$ alternatively and the procedures are summarized in Algorithm 4. Note that updating $\mathcal{P}_{6}$ does not decrease the objective function due to Algorithm 2, and updating $\mathcal{P}_{4}$ also does not decrease the objective function because Algorithm 3 leads to the global optimum. Hence, updating between $\mathcal{P}_{4}$ and $\mathcal{P}_{6}$ only leads to non-decreasing objective functions. It is trivial to see that the objective function of (23) is upper bounded and therefore Algorithm 4 is guaranteed to converge.
@@
0:  $\mathbf{P}^{\star}$ and $\mathbf{c}^{\star}$
@@
3:  repeat
4:     Update $R_{\text{c}}$, $R_{\text{p},k}$ based on $\mathbf{P}^{v}$.
5:     Update $\mathbf{c}$ using Algorithm 3.
6:     $\Delta\mathbf{P}:=g(R_{\text{max-min}})$ according to (28) and (29).
7:     Choose step size $t$ via Algorithm 2.
8:     Update. $\mathbf{P}^{v+1}:=\mathbf{P}^{v}+t\Delta\mathbf{P}$.
9:     Projection. $\mathbf{P}^{v+1}:=\frac{\sqrt{P_{\text{T}}}}{\|\mathbf{P}^{v+1}\|_{\text{F}}}\mathbf{P}^{v+1}$.
10:     Update. $R_{\text{max-min}}^{v+1}:=R_{\text{max-min}}(\mathbf{P}^{v+1})$.
11:     $v:=v+1$.
12:  until $\mid R_{\text{max-min}}^{v}-R_{\text{max-min}}^{v-1}\mid\;<\epsilon$ or $v\geq v_{\text{max}}$.
13:  return  $\mathbf{P}^{\star}:=\mathbf{P}^{v}$, $\mathbf{c}^{\star}=\mathbf{c}$.

### v2

@@ III-B Achievable Rate of RSMA
$+\frac{1}{|\mathcal{X}_{\mathcal{K}}|}\sum_{m=1}^{|\mathcal{X}_{\mathcal{K}}|}\log_{2}\sum_{l=1}^{|\mathcal{X}_{\mathcal{K}}|}\exp\Big{(}-\frac{|\mathbf{d}_{m,l}(\sum_{a\in\mathcal{K}}\mathbf{h}^{H}\mathbf{p}_{a}\mathcal{X}_{a})|^{2}}{2\sigma^{2}}\Big{)}$ (14)
@@
$+\frac{1}{|\mathcal{X}_{\mathcal{K}/k}|}\sum_{m=1}^{|\mathcal{X}_{\mathcal{K}/k}|}\log_{2}\sum_{l=1}^{|\mathcal{X}_{\mathcal{K}/k}|}\exp\Big{(}-\frac{|\mathbf{d}_{m,l}(\sum_{a\in\mathcal{K}/k}\mathbf{h}^{H}\mathbf{p}_{a}\mathcal{X}_{a})|^{2}}{2\sigma^{2}}\Big{)}$ (15)
$R_{\text{p},k}^{\text{SIC-free, approx.}}=$ $\log_{2}|\mathcal{X}_{k}|-\frac{1}{|\mathcal{X}_{\text{c}}\times\mathcal{X}_{\mathcal{K}}|}\sum_{m=1}^{|\mathcal{X}_{\text{c}}\times\mathcal{X}_{\mathcal{K}}|}\log_{2}\sum_{l=1}^{|\mathcal{X}_{\text{c}}\times\mathcal{X}_{\mathcal{K}}|}\exp\Big{(}-\frac{|\mathbf{d}_{m,l}(\mathbf{h}^{H}\mathbf{p}_{\text{c}}\mathcal{X}_{\text{c}}+\sum_{a\in\mathcal{K}}\mathbf{h}^{H}\mathbf{p}_{a}\mathcal{X}_{a})|^{2}}{2\sigma^{2}}\Big{)}$
@@
Because $s_{\text{c}}$ carries common part messages for potentially multiple users, $R_{\text{c}}$ is distributed among these users, and this distribution depends on the common part message each user contributed to $W_{\text{c}}$. Let $C_{k}$ denote the portion of the common stream rate that is allocated to user-$k$. Inherently, $C_{k}\geq 0$, $\forall k$, and $\sum_{k}C_{k}=R_{\text{c}}$. The rate experienced by user-$k$ is given by [^3] Hereinafter, the superscript ”SIC/SIC-free” indicates that the expression depends on whether SIC receivers or SIC-free receivers are used.
$R_{k}^{\text{SIC/SIC-free}}=C_{k}+R_{\text{p},k}^{\text{SIC/SIC-free}}.$ (17)
The performance loss incurred by removing SIC receiver at user-$k$ can be characterized as
$\begin{split}&R_{k}^{\text{SIC}}-R_{k}^{\text{SIC-free}}\\ =&I(s_{k};y_{k})-I(s_{k};y_{k}|s_{\text{c}})\\ =&I(s_{k};\sum_{k^{\prime}\in\mathcal{K}}\mathbf{h}_{k}^{H}\mathbf{p}_{k^{\prime}}s_{k^{\prime}}+\mathbf{h}_{k}^{H}\mathbf{p}_{\text{c}}s_{\text{c}}+n_{k})\\ &-I(s_{k};\sum_{k^{\prime}\in\mathcal{K}}\mathbf{h}_{k}^{H}\mathbf{p}_{k^{\prime}}s_{k^{\prime}}+n_{k}).\end{split}$ (18)
Understanding this difference requires addressing the following question: Consider a channel that takes $\mathbf{h}_{k}^{H}\mathbf{p}_{k}s_{k}$ as input and is disturbed by interference and noise denoted by $\sum_{k^{\prime}\in\mathcal{K}/k}\mathbf{h}_{k}^{H}\mathbf{p}_{k^{\prime}}s_{k^{\prime}}+n_{k}$. How detrimental is it to the achievable rate if an additional interference, $\mathbf{h}_{k}^{H}\mathbf{p}_{\text{c}}s_{\text{c}}$, is added? It is intuitive that $s_{\text{c}}$ is less detrimental to the achievable rate if it is from finite alphabet rather than Gaussian distributed, which motivated our analysis of SIC-free RSMA under finite alphabets. A recent work [37] addresses the impact of finite alphabet interference and shows that this impact is non-monotonic w.r.t. the interference power and can be low even if the interference is strong. Specifically, [37] shows that
$\begin{split}&\lim_{\mathrm{E}[|\mathbf{h}_{k}^{H}\mathbf{p}_{\text{c}}|^{2}]\rightarrow+\infty}I(s_{k};\sum_{k^{\prime}\in\mathcal{K}}\mathbf{h}_{k}^{H}\mathbf{p}_{k^{\prime}}s_{k^{\prime}}+\mathbf{h}_{k}^{H}\mathbf{p}_{\text{c}}s_{\text{c}}+n_{k})\\ =&I(s_{k};\sum_{k^{\prime}\in\mathcal{K}}\mathbf{h}_{k}^{H}\mathbf{p}_{k^{\prime}}s_{k^{\prime}}+n_{k}).\end{split}$ (19)
This indicates that the achievable rate of SIC-free RSMA can be made similar to SIC-based RSMA through proper precoding design.
@@ IV Precoding Optimization
In this section, we consider WSR and MMF maximization problems for RSMA with SIC and SIC-free receivers, and propose precoding optimization algorithms to address them.
@@ IV-A Weighted Sum-Rate Optimization
$\begin{split}\mathcal{P}_{1}:\;\;\underset{\mathbf{P},\mathbf{c}}{\max}\>\>&\sum_{k\in\mathcal{K}}u_{k}R_{k}^{\text{SIC/SIC-free}}\\ \text{s.t.}\>\>\>\;&\|\mathbf{P}\|_{\text{F}}^{2}\leq P_{\text{T}}\\ &\sum_{k\in\mathcal{K}}C_{k}=R_{\text{c}}\\ &\mathbf{c}\succcurlyeq\mathbf{0}^{1\times K},\end{split}$ (20)
@@
As a direct consequence of Proposition 1, one solution of $\mathbf{c}$ can already be determined given $u_{k}$, $\forall k\in\mathcal{K}$. As a result, $\mathbf{c}$ can be removed from the design variables, and $\mathcal{P}_{1}$ is simplified into
$\begin{split}\mathcal{P}_{2}:\;\;\underset{\mathbf{P}}{\max}\>\>&u_{k^{\prime}}R_{\text{c}}+\sum_{k\in\mathcal{K}}u_{k^{\prime}}R_{\text{p},k}^{\text{SIC/SIC-free}}\\ \text{s.t.}\>\>\>\;&\|\mathbf{P}\|_{\text{F}}^{2}\leq P_{\text{T}},\end{split}$ (21)
@@
The gradient of (6) is given by (22),
$\nabla_{\mathbf{P}}H(\mathbf{s}|y,\mathbf{h}^{H}\mathbf{P})^{\text{approx.}}=\frac{\log_{2}e}{|\mathcal{X}|}\sum_{m=1}^{|\mathcal{X}|}\frac{1}{\sum_{l=1}^{|\mathcal{X}|}u(\mathbf{P})}\sum_{l=1}^{|\mathcal{X}|}u(\mathbf{P})\left(-\frac{\mathbf{hh}^{H}\mathbf{P}\mathbf{d}_{m,l}(\mathcal{X})\mathbf{d}_{m,l}(\mathcal{X})^{H}}{2\sigma^{2}}\right)$ (22)
@@
$u(\mathbf{P})=\sum_{l=1}^{|\mathcal{X}|}\exp\Big{(}-\frac{|\mathbf{h}^{H}\mathbf{P}\mathbf{d}_{m,l}(\mathcal{X})|^{2}}{2\sigma^{2}}\Big{)}.$ (23)
Based on (22), $\nabla_{\mathbf{P}}R_{\text{c},k}$, $\nabla_{\mathbf{P}}R_{k}^{\text{SIC}}$ and $\nabla_{\mathbf{P}}R_{k}^{\text{SIC-free}}$ can be obtained trivially. One subgradient [^4] In this work, we use Clarke subdifferential for non-convex non-smooth functions [40]. of the WSR can be obtained as
$g(R_{\text{WSR}}^{\text{SIC/SIC-free}})=u_{k^{\prime}}\nabla_{\mathbf{P}}R_{\text{c},k^{\circ}}+\sum_{k\in\mathcal{K}}u_{k}\nabla_{\mathbf{P}}R_{k}^{\text{SIC/SIC-free}},$ (24)
where $k^{\circ}=\arg\min_{k\in\mathcal{K}}R_{\text{c},k}$.
It is easy to see that any solution of $\mathcal{P}_{2}$, denoted by $\mathbf{P}^{\star}$, must satisfy $\|\mathbf{P}^{\star}\|_{\text{F}}^{2}=P_{\text{T}}$. We therefore use the projection of $\mathbf{P}$ in $\{\mathbf{P}\big{|}\|\mathbf{P}\|_{\text{F}}^{2}=P_{\text{T}}\}$ as the updated solution after every iteration. The proposed Projected SubGradient Ascent (PSGA) algorithm is summarized in Algorithm 1555In Algorithm 1, depending on the assumption about receivers, $R_{\text{WSR}}=R_{\text{WSR}}^{\text{SIC}}\text{ or }R_{\text{WSR}}^{\text{SIC-free}}$.. For choosing the step length, $t$, we use a modified backtracking line search specific for PSGA as described in Algorithm 2. Compared to the classical backtracking for gradient descent/ascent, the proposed backtracking method incorporates projection when evaluating the predicted objective value after updating, ensuring a non-decreasing sequence of objectives. The convergence of general projected gradient methods is discussed in [41], although the convergence rate of such methods for non-convex problems is generally difficult to analyze. Note that each iteration in Algorithm 2 leads to a non-decreasing variation of the objective function, and the objective function of (20) is upper-bounded. Therefore, Algorithm 1 is guaranteed to converge. The convergence is also numerically evaluated in Figure 5, where a monotonic increase in the objective function and the convergence can be observed.
@@ IV-B Max-Min Fairness Optimization
$\begin{split}\mathcal{P}_{3}:\;\;\underset{\mathbf{P},\mathbf{c}}{\max}\>\>&\min_{k\in\mathcal{K}}\>\>R_{k}^{\text{SIC/SIC-free}}\\ \text{s.t.}\>\>\>\;&\|\mathbf{P}\|_{\text{F}}^{2}\leq P_{\text{T}}\\ &\sum_{k\in\mathcal{K}}C_{k}=R_{\text{c}}\\ &\mathbf{c}\succcurlyeq\mathbf{0}^{1\times K}.\end{split}$ (25)
Note that $\mathcal{P}_{3}$ is also a non-smooth non-convex problem and hence is difficult to solve. To develop efficient algorithms for $\mathcal{P}_{3}$, we decouple it into the following two sub-problems,
$\begin{split}\mathcal{P}_{4}:\;\;\underset{\mathbf{c}}{\text{max}}\>\>&\min_{k\in\mathcal{K}}\>\>R_{k}(\mathbf{c};\mathbf{P})\\ \text{s.t.}\>\>\>\;&\sum_{k\in\mathcal{K}}C_{k}=R_{\text{c}}\\ &\mathbf{c}\succcurlyeq\mathbf{0}^{1\times K},\end{split}$ (26)
@@
where $R_{k}(\mathbf{c};\mathbf{P})$ indicates the case where $R_{k}^{\text{SIC/SIC-free}}$ is determined with $\mathbf{P}$ as a given parameter and $\mathbf{c}$ as a variable, whereas $R_{k}(\mathbf{P};\mathbf{c})$ indicates the case in which $\mathbf{c}$ is given and $\mathbf{P}$ is the variable. Hence, $\mathcal{P}_{4}$ optimizes $\mathbf{c}$ with a fixed $\mathbf{P}$ while $\mathcal{P}_{5}$ optimizes $\mathbf{P}$ with a given $\mathbf{c}$.
@@
4:     $\mathbf{d}:=[R_{\text{p},1},...,R_{\text{p},k^{\prime}}]$.
5:     $\eta:=\frac{R_{\text{c}}+\mathbf{d}\mathbf{1}^{k^{\prime}\times 1}}{k^{\prime}}$.
6:     $\mathbf{c}^{\prime}:=\eta\mathbf{1}^{k^{\prime}\times 1}-\mathbf{d}$.
@@
The gap between point-wise minimum and its log-sum-exp approximation is characterized by the following [42].
$00\leq\min_{1\leq i\leq L}f_{i}-\frac{1}{\gamma}\log\sum_{i=1}^{L}\exp(\gamma f_{i})\leq\frac{1}{\gamma}\log L,\;\;\;\gamma<$ 0 (29)
Let $R_{\text{max-min}}$ denote the objective function in (28). One subgradient of $R_{\text{max-min}}$ w.r.t. $\mathbf{P}$ is derived to be
$g(R_{\text{max-min}})=\frac{\sum_{k\in\mathcal{K}}\exp\big{(}\gamma R_{k}(\mathbf{P};\mathbf{c})\big{)}g\big{(}R_{k}(\mathbf{P};\mathbf{c})\big{)}}{\sum_{k\in\mathcal{K}}\exp\big{(}\gamma R_{k}(\mathbf{P};\mathbf{c})\big{)}},$ (30)
@@
$g\left(R_{k}(\mathbf{P};\mathbf{c})\right)=c_{k}\nabla_{\mathbf{P}}R_{\text{c},k^{\circ}}+\nabla_{\mathbf{P}}R_{\text{p},k},$ (31)
and $k^{\circ}=\arg\min_{k}R_{\text{c},k}$. Similarly to $\mathcal{P}_{2}$, $\mathcal{P}_{6}$ can also be solved using PSGA. The solution of (25) is obtained by updating the solution for $\mathcal{P}_{4}$ and $\mathcal{P}_{6}$ alternatively, and the procedures are summarized in Algorithm 4. Note that, in Algorithm 4, the updates on $\mathbf{P}$ do not decrease the objective function due to the step sizes selected by Algorithm 2, and the updates on $\mathbf{c}$ also do not decrease the objective function, because Algorithm 3 leads to the global optimum. Hence, every iteration in Algorithm 4 leads to a non-decreasing variation of the objective function. It is trivial to see that the objective function of $\mathcal{P}_{3}$ is upper-bounded. Therefore, Algorithm 4 is guaranteed to converge.
@@
0:  $\mathbf{P}^{\star}$ and $\mathbf{c}^{\star}$.
@@
3:  Update. $R_{\text{max-min}}^{0}:=R_{\text{max-min}}(\mathbf{P}^{0})$.
4:  repeat
5:     Update $R_{\text{c}}$, $R_{\text{p},k}$ based on $\mathbf{P}^{v}$.
6:     Update $\mathbf{c}$ using Algorithm 3.
7:     $\Delta\mathbf{P}:=g(R_{\text{max-min}})$ according to (30) and (31).
8:     Choose step size $t$ via Algorithm 2.
9:     Update. $\mathbf{P}^{v+1}:=\mathbf{P}^{v}+t\Delta\mathbf{P}$.
10:     Projection. $\mathbf{P}^{v+1}:=\frac{\sqrt{P_{\text{T}}}}{\|\mathbf{P}^{v+1}\|_{\text{F}}}\mathbf{P}^{v+1}$.
11:     Update. $R_{\text{max-min}}^{v+1}:=R_{\text{max-min}}(\mathbf{P}^{v+1})$.
12:     $v:=v+1$.
13:  until $\mid R_{\text{max-min}}^{v}-R_{\text{max-min}}^{v-1}\mid\;<\epsilon$ or $v\geq v_{\text{max}}$.
14:  return  $\mathbf{P}^{\star}:=\mathbf{P}^{v}$, $\mathbf{c}^{\star}=\mathbf{c}$.